# Decoding Speculative Decoding

### v3

@@ 4 Draft Model Design for Speculative Decoding
The above results indicate that to improve the throughput of speculative decoding, it is necessary to improve the latency of draft models, i.e., can we design a model that provides a similar TAR at a lower inference cost? In the next section, we study the possibility of such a design based on the above insights.
4.1 Draft Model Design
Method: We leverage recent advances in structured LLM pruning, Sheared-LLaMa Xia et al. (2023b), which provides a framework to prune larger models to a specified smaller configuration. Sheared-LLaMa Xia et al. (2023b) learns layers, attention heads, and neurons to mask from the large model to prune it into the specified small model. The flexibility enables us to prune LLaMa-7B into desirable model configurations. In our experiments, we prune our models from LLaMa-7B using 0.4B tokens sampled from the RedPajama Dataset Computer (2023) (the same as in Sheared-LLaMa Xia et al. (2023b)), but skipped the expensive fine-tuning step on 50B more tokens (and hence the name NoFT). We find that this is sufficient to achieve a significantly higher throughput.
Deep vs Wide Model Comparison: Our goal is to start with LLaMa-7B and produce a wider version of Sheared-LLaMa-1.3B while keeping the number of parameters the same as in Sheared-LLaMa-1.3B. We choose Sheared-LLaMa-1.3B since it achieves the highest throughput in our benchmark among existing models (Blue dots in Figure 6). We use two configurations: the first configuration was provided by the Sheared-LLaMa authors (NoFT-1.3B), and we designed the second configuration (NoFT-Wide-1.3B) to optimize for better speculative decoding throughput. Table 1 shows the detailed configuration of the two models. We slash the number of layers by half, from 24 to 12, and keep the total parameter count roughly the same by increasing the intermediate size from 5504 to 9280, the number of attention heads from 16 to 20, and the corresponding model dimension from 2048 to 2560.
Figure 6(a)6(b), and 8(b) show that we can achieve up to $30\%$ higher speculative decoding throughput using only $0.8\%$ of tokens used to train Sheared-LLaMa-1.3B.
Table 1: This table shows the model configuration of the two pruned models. Here l represents the number of layers, h represents the number of attention heads, $d_{\text{inter}}$ represents intermediate size, and $d_{\text{model}}$ represents model dimension.
@@ 4.1 Draft Model Design
Furthuer, our model also reduces intermediate memory usage by $37\%$ compared to the original model. Our memory saving mainly comes from wider MLP layers, which does not require storing KV-caches.
[…]
Table 2: This table shows the speculative decoding throughput and the latencies to generate 8 tokens using the two pruned draft models.
Draft Model TAR Latency (ms) Tput (tokens/s)
NoFT-1.3B 3.81 105.1 23.10
NoFT-Wide-1.3B 3.70 53.5 32.59
4.2 Ablation Studies
Temperature sampling: Next, we perform ablation by varying the sampling method. In our prior experiments, we used the greedy decoding method. However, sampling-based decoding methods often improve LLM decoding quality Holtzman et al. (2019); Ficler and Goldberg (2017); Fan et al. (2018). To understand the effect of sampling on our findings, we performed experiments with temperature sampling by choosing temperatures of 0.5 and 1. In Table 3, we show that our model achieves even higher speedup than existing models when stochasticity is introduced to the decoding process. On three datasets we achieved from $97\%$ to $111\%$ higher throughput using our NoFT-Wide-796M model compared to using an existing fine-tuned model. The speedups in different temperature settings show that our speedups generalize across different sampling methods.
Table 3: This table shows the throughput of speculative decoding (tokens/s) with temperature sampling on various datasets and models. Our NoFT-Wide-796M model achieved $97.7\%$ to $111.8\%$ higher throughput comparing to the existing Sheared-LLaMA-1.3B model while only using $0.8\%$ of its training tokens.
@@ 4.2 Ablation Studies
Varying the target model: Prior experiments are performed with LLaMa-65B as the target model. As newer generations of models roll out, we would like to see if our conclusion holds on newer generations of models. In this ablation study, we evaluate our best NoFT-Wide-796M model against LLaMa-2-70B model. Table 4 shows that though our NoFT-Wide-796M is distilled from LLaMa-7B, it can achieve a similar token acceptance rate when the target model is from LLaMa-2 family. We believe the similarity in performance is due to similar training datasets. This shows that our distilled model can be applied to future families of models based on similar training recipes with little to no changes.
@@
LLaMa-2 70B 2.55 2.68 2.64
Supervised fine-tuned models: Prior experiments are performed on base models to study the scaling of draft models. In practice, supervised fine-tuned models are adopted for their better instruction-following capabilities. In this section, we compare our best NoFT-Wide-796M model to Tiny-LLaMa-1.1B with Vicuna 33B as the target model. Note that our NoFT-Wide-796M is pruned from the base version of LLaMa-7B without fine-tuning. Table 5 shows that NoFT-Wide-796M outperforms Tiny-LLaMa-1.1B in all cases by up to $45\%$. While Tiny-LLaMa-1.1B has a TAR $35\%$ and $32\%$ higher than NoFT-Wide-796M on MMLU and Hellaswag, respectively, its latency is 4x higher due to having 22 layers in the model compared to NoFT-Wide-796M with merely 5 layers. This ablation study also demonstrates how speculative decoding is bottlenecked by draft model depth and that a draft model obtained from the non-fine-tuned base model, when appropriately designed, can still provide significant speedup over draft models fine-tuned for chatbot purposes.
@@
5 Discussion
Next, we discuss how our insights can change if the models or the underlying hardware change.
5.1 Future Draft Model Design
@@
$Tput=\begin{cases}\displaystyle\frac{TAR}{(t^{d}_{target}+t^{d}_{draft})}&\text{if }TAR>1,\\ \displaystyle\frac{1}{(t^{d}_{target}+t^{d}_{draft})}&\text{if }TAR\leq 1.\end{cases}\vspace{-3pt}$
In the Appendix (Figure 12), we show that this simplified formula almost perfectly captures the real speculative decoding throughput. Here, $t^{d}$ represents the latency to generate $d$ tokens autoregressively. In this section, with the aid of the performance model, we provide quantitative answers to several questions: First, we study the improvement in TAR a larger draft model needs to be provided to compensate for the additional inference cost. Next, we study how much improvement in latency is required to change the choice of the draft model.
@@
Improvement in TAR needed to switch to a larger draft model model.
In Figure 2, we observed that with existing datasets and models, we are better off with the smallest model as the draft model, e.g., OPT-125M, than choosing a larger model. However, there is a possibility that the TAR difference will become greater for new datasets. In Figure 7, we plot the improvement in TAR (extra TAR), which larger models in the OPT model family should provide to match the throughput of the smallest model (OPT-125M) for MMLU. We find that if a 1.3B model can achieve a TAR advantage greater than 2 over OPT-125M for a new workload, we would choose the 1.3B model instead. Furthermore, given that the maximum TAR is capped at 8 in our scenario due to the length of draft token generation, it becomes unfeasible for OPT-2.7B and OPT-6.7B to surpass OPT-125M in performance. This is because the improvement needed in TAR for OPT-6.7B to match the throughput of OPT-125M would exceed this maximum limit.
Improvement in latency for switching to higher TAR model.
As hardware evolves, latency scaling patterns may change with more computing power and memory bandwidth. Therefore, conclusions drawn on specific hardware (e.g., A100) may not hold for newer or older hardware (e.g., H100 or V100). To account for changing hardware, we study how much draft model latency improvement is needed to achieve throughput parity. To demonstrate this, we first compute the latency reduction needed for different members in OPT family to reach the same throughput as the smallest draft model in Table 6. We find that up to $56\%$ of latency reduction is needed to achieve the same throughput. For instance, for OPT-1.3B to achieve parity throughput with OPT-125M, its latency needs to be reduced by 32.9%. This reinforces our finding that latency reduction provided by the smaller models has significantly more benefit than the extra TAR provided by a larger draft model.
6 Conclusion
[…]
7 Limitations
Our work aims to improve the inference efficiency of LLMs by designing better draft models for speculative decoding. Since speculative decoding preserves the output from the LLM, our work will not amplify existing biases in LLMs. However, limiting and reducing such biases are out of the scope of this work. Furthermore, since we are making LLM generation more efficient, we believe our work will not have a significant negative environmental impact.
Appendix A More Experiment Results
Figure 9: Performance Breakdown of LLaMa speculative decoding, lookahead length is set to be optimal lookahead length found empirically.
(a) Model accuracy vs. TAR for OPT models on MMLU and Hellaswag
(b) Model accuracy vs TAR for LLaMa models on MMLU and Hellaswag
Required TAR to match throughput.
We can also use our analytical model to predict the TAR necessary for different models to achieve a target throughput. This can be useful in scenarios where developers deploy speculative decoding-based LLMs and must meet a throughput goal. In Figure 11, we plot the TAR needed by existing models to achieve a specific throughput.
@@ Appendix A More Experiment Results
Appendix B OPT-350M Configurations
@@
Table 7: This table shows model configuration of various OPT-350M models we created. The goal is to explore the tradeoff between model depth and width while keeping the total parameter count constant.
@@
Table 8: This table shows the latency of each auto-regressive generation step of the draft model.
Model Latency (ms)
OPT-125M 6.23
OPT-350M 11.74
OPT-1.3B 12.64
OPT-2.7B 16.35
OPT-6.7B 18.56
Appendix C Simplifying Analytical model
@@
$Tput=\begin{cases}\displaystyle\frac{TAR}{(t^{d}_{target}+t^{d}_{draft})}&\text{if }TAR>1,\\ \displaystyle\frac{1}{(t^{d}_{target}+t^{d}_{draft})}&\text{if }TAR\leq 1.\end{cases}$
@@
Verifying Analytical Model
In Figure 12, we compare the throughput predicted by our model with throughput measured on real hardware for two model families: LLaMa (7B and 13B) and OPT (125M, 350M, 1.3B, 2.7B, and 6.7B) to serve LLaMa-65B and OPT-66B on MMLU.

### v4

@@ 4 Draft Model Design for Speculative Decoding
The above results indicate that to improve the throughput of speculative decoding, it is necessary to improve the latency of draft models, i.e., can we design a model that provides a similar TAR at a lower inference cost? In the next section, we study the possibility of such a design.
Table 1: This table shows the model configuration of the two pruned models. Here l represents the number of layers, h is the number of attention heads, $d_{\text{inter}}$ is intermediate size, and $d_{\text{model}}$ is model dimension.
@@
(a) LLaMa Series on MMLU
(b) LLaMa Series on Hellaswag
Table 2: This table shows the throughput of speculative decoding (tokens/s) with temperature sampling on various datasets and models. Our NoFT-Wide-796M model achieved $97.7\%$ to $111.8\%$ higher throughput compared to the existing Sheared-LLaMA-1.3B model while only using $0.8\%$ of its training tokens.
@@
4.1 Draft Model Design
Method: We leverage recent advances in structured LLM pruning, Sheared-LLaMa Xia et al. (2023b), which provides a framework to prune larger models to a specified smaller configuration. Sheared-LLaMa Xia et al. (2023b) learns layers, attention heads, and neurons to mask from the large model to prune it into the specified small model. The flexibility enables us to prune LLaMa-7B into desirable model configurations. In our experiments, we pruned our models from LLaMa-7B using 0.4B tokens sampled from the RedPajama Dataset Computer (2023) following Xia et al. (2023b) but skipped the expensive fine-tuning step on 50B more tokens (and hence the name NoFT). We find that this is sufficient to achieve a significantly higher throughput.
Table 3: This table shows the TAR, per-step decoding latency, and speculative decoding throughput using the two pruned draft models.
Draft Model TAR Latency Throughput
(ms) (tokens/s)
NoFT-1.3B 3.81 13.13 23.10
NoFT-Wide-1.3B 3.70 6.69 32.59
Deep vs wide model comparison: Our goal is to start with LLaMa-7B and produce a wider version of Sheared-LLaMa-1.3B while keeping the number of parameters the same as in Sheared-LLaMa-1.3B. We choose Sheared-LLaMa-1.3B since it achieves the highest throughput in our benchmark among existing models (blue dots in Figure 6). We use two configurations: the first configuration was provided by the Sheared-LLaMa authors (NoFT-1.3B), and we designed the second configuration (NoFT-Wide-1.3B) to optimize for better speculative decoding throughput. Table 1 shows the detailed configuration of the two models. We slash the number of layers by half, from 24 to 12, and keep the total parameter count roughly the same by increasing the intermediate size from 5504 to 9280, the number of attention heads from 16 to 20, and the corresponding model dimension from 2048 to 2560. Figure 6(a)6(b), and 7(b) show that we can achieve up to $30\%$ higher speculative decoding throughput using only $0.8\%$ of tokens used to train Sheared-LLaMa-1.3B.
[…]
We also verified our model’s robustness under different temperatures. We performed experiments with temperature sampling by choosing temperatures of 0.5 and 1 following Leviathan et al. (2023). In Table 2, we show that our model achieves even higher speedup than existing models when stochasticity is introduced to the decoding process. On three datasets we achieved from $97\%$ to $111\%$ higher throughput using our NoFT-Wide-796M model compared to using an existing fine-tuned model. The speedups in different temperature settings show that our speedups generalize across different sampling methods.
Table 4: This table shows the memory usage of our draft model LLaMa-Wide-1.3B versus Sheared-LLaMa-1.3B. Our draft model reduces KV cache by $37\%$.
KV cache Activations
Context Length Sheared Wide Sheared Wide
256 48MB 30MB 1MB 1.25MB
512 96MB 60MB 2MB 2.5MB
1024 192MB 120MB 4MB 5MB
Table 5: This table shows the prefill and per-step autoregressive decoding latency for different batch sizes.
Batch size Prefill (s) Autoregressive decoding (s)
Model LLaMA-65B LLaMA-NoFT Sheared
Wide-1.3B LLaMA-1.3B
1 0.060 0.0097 0.018
4 0.061 0.0097 0.019
16 0.063 0.0099 0.019
32 0.065 0.0101 0.019
Batch size and KV cache: By designing a shallower model with wider MLP layers, we reduce the KV-cache size required by the model. This allows us to increase the batch size we can accommodate. Table 4 shows that for the same sized 1.3B model, our wider design reduces KV-Cache by more than $37\%$. Table 5 demonstrates that increasing the batch size has minimal impact on autoregressive decoding latency, thereby increasing throughput.
LLaMa-3 Results We apply our approach to the newest LLaMa-3.2-1B model. All other experiment settings follow section 4.1. We evaluated on MMLU, Hellaswag, and Chatbot Arena prompts. We compare our Wide-829M model against LLaMa-3.2 and self-speculative decoding. Since LLaMa-3 is trained on more tokens than LLaMa-2 (15 trillion vs 2 trillion), pruning its layers had a larger impact on TAR. We follow prior work Muralidharan et al. (2024) to recover the TAR after pruning by distilling it over 1 million tokens. Note that this process can be finished within 10 minutes. In this experiment, we use LLaMa-3.1-8B as the target model. Table 6 shows that our pruned Wide-829M model achieves $42.6\%$ higher throughput compared to LLaMa-3.2-1B and $51.8\%$ higher throughput compared to self-speculative decoding. This experiment shows that our approach works well on SoTA models.
Table 6: This table shows the speculative decoding throughput and the latencies of our Wide-829M models against LLaMa-3.2-1B model and self-speculative decoding.
Throughput(tokens/s)
Draft Model Chat MMLU Hellaswag
Wide-829M 53.35 61.77 60.44
LLaMA-3.2-1B 37.41 44.89 39.03
Self-Speculative 35.14 40.41 40.02
4.2 Ablation Studies
@@
Varying the target model: Prior experiments are performed with LLaMa-65B as the target model. As newer generations of models roll out, we would like to see if our conclusion holds on newer generations of models. In this ablation study, we evaluate our best NoFT-Wide-796M model against the LLaMa-2-70B model. Table 7 shows that though our NoFT-Wide-796M is distilled from LLaMa-7B, it can achieve a similar token acceptance rate when the target model is from LLaMa-2 family. This shows that our distilled model can be applied to various models based on similar training recipes and tokenizers. We also show more results on the newest LLaMa-3.1 and LLaMa-3.2 families in Appendix 5.
@@
LLaMa-2-70B 2.55 2.68 2.64
Supervised fine-tuned models: Prior experiments are performed on base models to study the scaling of draft models. In practice, supervised fine-tuned models are adopted for their better instruction-following capabilities. In this section, we compare our best NoFT-Wide-796M model to Tiny-LLaMa-1.1B with Vicuna 33B as the target model. Note that our NoFT-Wide-796M is pruned from the base version of LLaMa-7B without fine-tuning. Table 8 shows that NoFT-Wide-796M outperforms Tiny-LLaMa-1.1B in all cases by up to $45\%$. While Tiny-LLaMa-1.1B has a TAR $35\%$ and $32\%$ higher than NoFT-Wide-796M on MMLU and Hellaswag, respectively, its latency is 4x higher due to having 22 layers in the model compared to NoFT-Wide-796M with merely 5 layers. This ablation study also demonstrates how speculative decoding is bottlenecked by draft model depth and that a draft model obtained from the non-fine-tuned base model, when appropriately designed, can outperform fine-tuned models.
@@
Comparison against self-speculative decoding: Self-speculative decoding Zhang et al. (2023) was proposed to leverage the base model to avoid the need to train a draft model. However, we show that we achieve significantly higher throughput with our proposed model architecture design. We pick the settings where self-speculative decoding achieves its highest speedup (LLaMa-2-70B as target model with greedy sampling on CNNDM and XSum datasets). However, our LLaMa-NoFT-Wide-796M outperforms self-speculative decoding by up to $53\%$. This is because our LLaMa-NoFT-Wide-796M model’s autoregressive decoding latency is only $7\%$ of the target 70B model’s prefill latency. Self-speculative decoding suffers from a significant drop in drafting accuracy after dropping more than 42 layers out of 80 in LLaMa-2-13B Zhang et al. (2023). Table 10 shows that on a 70B target model, they would still need to compute 33 attention layers and 43 MLP layers to obtain a draft response, while our LLaMa-NoFT-Wide-796M only has 5 attention and MLP layers each. Therefore, we obtain a much higher throughput compared to self-speculative decoding. More comparisons on LLaMa-3 are presented in Appendix 5.
Table 9: This table shows the number of layers used in self-speculative decoding and our draft model.
Layers
Model Setup Attention MLP
Self-Speculative LLaMa-2-13B 16 30
LLaMa-2-70B 33 43
Draft Model NoFT-Wide-796M 5 5
Table 10: This table shows the speedup achieved by our LLaMa-NoFT-Wide-796M and self-speculative decoding, respectively.
Strategy Speedup
CNNDM XSum
Self-Speculative 99% 60%
NoFT-Wide-796M 131% 145%
5 Conclusion
[…]
Limitations
Our work aims to improve the inference efficiency of LLMs by designing better draft models for speculative decoding. One limitation of our work is that we focus on empirically studying the performance bottleneck and improving lossless speculative decoding throughput, where lossless refers to preserving the target LLM’s output distribution. Broader studies on approximate efficient inference algorithms are left for future work.
Acknowledgement
We gratefully acknowledge the support of the NSF Diamond project OAC-2311767 (Democratizing Large Neural Network Model Training for Science). We thank Kangwook Lee for his helpful suggestions during the project. This research used resources from the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility, using the NERSC award DDR-ERCAP0029980. This research also used computational resources from the NSF Cloudlab Duplyakin et al. (2019) facility.
Appendix A More Experiment Results
Table 11: This table shows the latency of each autoregressive generation step of the draft model.
Model Latency (ms)
OPT-125M 6.23
OPT-350M 11.74
OPT-1.3B 12.64
OPT-2.7B 16.35
OPT-6.7B 18.56
A.1 OPT analysis
(a) Model accuracy vs TAR for OPT models
(b) Model accuracy vs TAR for LLaMa models
Figure 9: This figure shows the performance breakdown of LLaMa speculative decoding, look ahead length is set to be optimal look ahead length found empirically.
Appendix B Discussion
In this section, we discuss how our insights can change if the models or the underlying hardware change.
B.1 Future Draft Model Design
@@
$\text{Throughput}=\begin{cases}\displaystyle\frac{TAR}{(t^{d}_{target}+t^{d}_{draft})}&\text{if }TAR>1,\\ \displaystyle\frac{1}{(t^{d}_{target}+t^{d}_{draft})}&\text{if }TAR\leq 1.\end{cases}\vspace{-3pt}$
We show that this simplified formula almost perfectly captures the real speculative decoding throughput. Here, $t^{d}$ represents the latency to generate $d$ tokens autoregressively. In this section, with the aid of the performance model, we provide quantitative answers to several questions: First, we study the improvement in TAR a larger draft model needs to be provided to compensate for the additional inference cost. Next, we study how much improvement in latency is required to change the choice of the draft model.
@@
Required TAR to match throughput: We use our analytical model to predict the TAR necessary for different models to achieve a target throughput. This can be useful in scenarios where developers deploy speculative decoding-based LLMs and must meet a throughput goal. In Figure 10, we plot the TAR needed by existing models to achieve a specific throughput.
@@
Improvement in TAR needed to switch to a larger draft model: In Figure 2, we observed that with existing datasets and models, we are better off with the smallest model as the draft model, e.g., OPT-125M, than choosing a larger model. However, there is a possibility that the TAR difference will become greater for new datasets. In Figure 11, we plot the improvement in TAR (extra TAR), which larger models in the OPT model family should provide to match the throughput of the smallest model (OPT-125M) for MMLU. We find that if a 1.3B model can achieve a TAR advantage greater than 2 over OPT-125M for a new workload, we would choose the 1.3B model instead. Furthermore, given that the maximum TAR is capped at 8 in our scenario due to the length of draft token generation, it becomes unfeasible for OPT-2.7B and OPT-6.7B to surpass OPT-125M in performance. This is because the improvement needed in TAR for OPT-6.7B to match the throughput of OPT-125M would exceed this maximum limit.
Improvement in latency for switching to higher TAR model: As hardware evolves, latency scaling patterns may change with more computing power and memory bandwidth. Therefore, conclusions drawn on specific hardware (e.g., A100) may not hold for newer or older hardware (e.g., H100 or V100). To account for changing hardware, we study how much draft model latency improvement is needed to achieve throughput parity. To demonstrate this, we first compute the latency reduction needed for different members in OPT family to reach the same throughput as the smallest draft model in Table 12. We find that up to $56\%$ of latency reduction is needed to achieve the same throughput. For instance, for OPT-1.3B to achieve parity throughput with OPT-125M, its latency needs to be reduced by 32.9%. This reinforces our finding that latency reduction provided by the smaller models has significantly more benefit than the extra TAR provided by a larger draft model.
Appendix C OPT-350M Configurations
@@
Table 13: This table shows the model configuration of various OPT-350M models we created. The goal is to explore the tradeoff between model depth and width while keeping the total parameter count constant.
@@
Appendix D Simplifying Analytical model
@@
$\text{Throughput}=\begin{cases}\displaystyle\frac{TAR}{(t^{d}_{target}+t^{d}_{draft})}&\text{if }TAR>1,\\ \displaystyle\frac{1}{(t^{d}_{target}+t^{d}_{draft})}&\text{if }TAR\leq 1.\end{cases}$
@@
Verifying analytical model: In Figure 12, we compare the throughput predicted by our model with throughput measured on real hardware for two model families: LLaMa (7B and 13B) and OPT (125M, 350M, 1.3B, 2.7B, and 6.7B) to serve LLaMa-65B and OPT-66B on MMLU.